### v2

# inkscapepath=inkscape, inkscapearea=page, clean=true, tex=false

Abstract
[…]
University of Zurich University of Zurich
Zurich, Switzerland Zurich, Switzerland
laura.wagner@khist.uzh.ch eva.cetinic@uzh.ch
Content warning: This paper contains references to pornography, gender-based digital violence and verbal vulgarity that might be disturbing, distressing, and/or offensive.
[…]
University of Zurich University of Zurich
Zurich, Switzerland Zurich, Switzerland
laura.wagner@khist.uzh.ch eva.cetinic@uzh.ch
Content warning: This paper contains references to pornography, gender-based digital violence and verbal vulgarity that might be disturbing, distressing, and/or offensive.
Keywords: Generative AI, Text-to-Image AI, Open-Source AI, AI Platformization, AI Democratization, Community-Driven AI, AI and Visual Culture, Ethical AI, Gender Bias in AI, NSFW Content in AI, CivitAI, Digital Feminism, Model Personalization
1 Introduction
The evolution of open-source text-to-image (TTI) technology has shifted from being a niche area for early adopters to becoming a vast ecosystem driven by increasing investment and commercialization. The technology continues to advance rapidly, with recent innovations allowing users to personalize models through model adapters, tailoring them for specific tasks and creative pursuits such as mimicking a specific art style or reproducing the physiognomy of a particular individual. While model personalizations open up unprecedented opportunities for artists and researchers, they also raise significant ethical and legal concerns, particularly the potential for misuse such as non-consensual deepfakes and disinformation. Platforms like CivitAI [^1] CivitAI https://civitai.com/, (Accessed: 11 November 2024), a pioneering and established hub for developing and sharing open-source TTI models, have played a central role in the expansion of AI-generated content and specifically TTI models.
What sets model-sharing hubs apart from other platforms is their function: the content shared on these hubs is not an end product but a tool for generating limitless derivative content. Personalized models, can act as powerful multipliers as they enable users to create vast amounts of content that mimics the characteristics of the original training data, which can then be widely shared across social media and other online platforms. This dynamic harbors a high potential for virulence, as its impact is expected to extend far beyond the confines of the communities creating these models. We are particularly concerned about misuse, such as technology-facilitated gender-based violence (TFGBV) (UN Women, 2024), which is facilitated by systems capable of creating entirely fabricated, photorealistic, and believable imagery within seconds. We aim to highlight how model-sharing platforms may exacerbate these concerning developments in their role of fostering environments that are tolerant of abusive behavior, while becoming increasingly hostile towards certain user groups. This trajectory, if left unchecked, may prove more detrimental to open-source innovation than any external regulation.
Drawing on a definition of misogyny as a set of hostile mechanisms designed to reinforce patriarchal power structures (Manne, 2017), our objective is to investigate how these dynamics manifest themselves in open-source TTI communities. Specifically, we intend to explore how over-sexualization, objectification, and the erosion of consent in TTI-generated content lead to the proliferation of harmful stereotypes and the marginalization of certain demographic groups, specifically women [^2] We use the terms women and girls to refer to female-presenting individuals. We recognize that gender presentation and identity are diverse, but use simplified terms here for clarity..
Figure 1: Random sample of images taken from CivitAI in March 2024 with different NSFW levels as assigned by the platform.
In practice, our study seeks to examine the underlying mechanisms contributing to the escalation of NSFW content and the perpetuation of misogynistic tropes within open-source TTI communities, as exemplified by the platform CivitAI. The dominance of NSFW content is illustrated in Figure 1, which shows a random sample of images taken from CivitAI in March 2024, as categorized by the platform according to different levels of NSFW.
In our study, we perform a detailed analysis of user-generated content, model adapters, and textual training data in order to uncover how community-driven, tool-driven, and model-infused biases may have intensified within the platform’s ecosystem over the last two years. Our results suggest a pattern of systematic misogyny that contributes to a self-reinforcing cycle of platform degradation, facilitating the spread and normalization of harmful, exclusionary, and exploitative content.
Besides acknowledging these issues, we identify potential areas for intervention, such as improving tools used for training models and model adapters, vetting model metadata, and developing and enforcing ethical frameworks. Through these efforts, our study aims to contribute to a more inclusive open-source TTI ecosystem.
2 Understanding Model Personalization and Model-Sharing Hubs: A Case Study of CivitAI
In early 2021, OpenAI demonstrated the capabilities of its novel TTI system, DALL·E, and showcased its capabilities with the now legendary depictions of avocado chairs (OpenAI, 2021), exemplifying the model’s ability to combine distinct concepts to produce novel outputs. OpenAI did not release DALL·E in full, yet made one of its components, the CLIP model (Contrastive Language-Image Pre-Training) (Radford et al., 2021), publicly available. CLIP was trained on a large dataset of image–text pairs from the internet, allowing it to associate visuals with descriptive language in a highly meaningful way. This breakthrough transformed multimodal machine learning by enabling direct comparison between images and text, a capability later used in generative models like DALL·E to produce semantically accurate images from prompts.
The open-source community sought to replicate DALL·E’s functionality by employing CLIP for various implementations of generative image synthesis, such as CLIP + VQGAN (Crowson et al., 2022) or CLIP-Guided Diffusion (Crowson, 2021). The first implementations of community-developed TTI pipelines were often shared via social media and Discord channels among open-source enthusiasts. These early communities and first implementations were characterized by collaborative efforts, with a focus on creative and artistic applications (Oppenlaender, 2022).
In August 2022, the release of Stable Diffusion by CompVis, Runway AI, and Stability AI (Rombach et al., 2022) marked a significant milestone as the first open-source TTI foundation model. Its popularity due to high-quality image generation capabilities indicated a shift from community-driven innovation toward investor-funded commercial activity. Concurrently, increased scrutiny arose concerning copyright violations, attribution challenges, and social biases resulting from training these models on uncurated, web-scraped datasets. Systematic research efforts emerged to quantify and evaluate gender, racial, and cultural biases in generated images (Luccioni et al., 2023; Wu et al., 2024; Naik and Nushi, 2023; Zhang et al., 2024b; Wang et al., 2024; Ghosh and Caliskan, 2023; Kannen et al., 2024; Cetinic, 2022; Zhang et al., 2024a).
New developments in TTI technology, notably model personalization through adapters such as LoRA (Low-Rank Adaptation) (Hu et al., 2022), originally developed for language models and later adapted for TTI, and textual inversion (Gal et al., 2023), offered refined control over generated content without retraining entire models. Applications with graphical user interfaces, such as Stable-Diffusion-WebUI (automatic1111, 2022) and ComfyUI (comfyanonymous, 2023), lowered the barrier to entry for TTI, allowing users to personalize models easily and without advanced technical skills. However, model personalization introduced new risks, particularly through the tailored generation of deepfake imagery, exacerbating digital sexual violence disproportionately affecting women (Viola and Voto, 2023). Accessibility increased significantly through model-sharing platforms such as Hugging Face, Replicate, and notably CivitAI, intensifying these potential harms.
Despite the prominence of these platforms, research into their implications remains limited. Existing studies identified prevalent NSFW content, abusive generative practices, and sexually exploitative AI-generated imagery on platforms like CivitAI (Wei et al., 2024; Palmini et al., 2024; Maiberg, 2024, 2023). Broader analyses indicate challenges in governance, highlighting platforms’ antagonistic stance toward artists’ rights and inadequacies of traditional moderation frameworks in addressing these issues (Gorwa and Veale, 2024; Cui and Araujo, 2024).
Our exploratory analysis of CivitAI aims to quantify and contextualize these systemic challenges, emphasizing community dynamics and misogynistic elements. Understanding the layered structure through which content is produced and shaped, including generative models and the surrounding sociotechnical ecosystem, offers valuable insight and supports a more holistic perspective. Figure 2 provides a schematic overview of the CivitAI platform, illustrating the layers of content generation and circulation.
Figure 2: Schematic overview of the CivitAI platform, showing layered relationships between the curated front-end images, the underlying generative models, and the broader sociotechnical ecosystem. This includes tools, data sources, and cultural influences, that shape model development.
2.1 Data Collection and Categorization
To illustrate the evolving dynamics of this prominent model-sharing hub, we collected and analyzed image metadata, model metadata, and training process metadata from content published on CivitAI over a 2-year period (late November 2022 to December 2024), using the CivitAI REST API (CivitAI, 2023). The dataset covers 40,630,560 images and 231,252 models as of 19th January [^3] Access available upon request: https://huggingface.co/datasets/pm-paper-datasets. The model metadata includes 9,948 standalone models, commonly referred to as “checkpoints”, capable of generating images independently, and 221,304 adapters (e.g., LoRA, Textual Inversion) that require a checkpoint to function. We analyzed three primary data categories: image metadata, model metadata, and training process metadata.
Image metadata included content organization metadata and generation metadata. Content organization metadata covered identifiers (CivitAI ID, perceptual hash), creator usernames, timestamps, engagement metrics (likes), and NSFW categories. CivitAI classifies images into five NSFW levels: Level 1 (PG) (SFW); Level 2 (PG-13) (revealing clothing, violence); Level 4 (R) (adult themes, partial nudity); Level 8 (X) (graphic nudity); and Level 16 (XXX) (explicit sexual or disturbing content) using Amazon Rekognition and manual review (CivitAI, 2025). Additionally, the image metadata encompassed generation settings such as “configuration scale” and “random seed”, as well as models and adapters used (represented by CivitAI IDs or model hashes). For an example structure, see Figure 10(a) in the Supplementary Material.
Model-related metadata included unique CivitAI IDs, version-specific IDs, download counts, timestamps, engagement metrics, a POI (person of interest) flag indicating mimicry of real individuals. Metadata from model creation pipelines, including base model details and framework-specific hashes, facilitated identification. A detailed structure is shown in Figure 10(b) in the Supplementary Material.
Finally, we extracted metadata written during training of a model, which were used to identify the auto-tagging system responsible for annotating the training data for model personalizations. This metadata was extracted from the downloaded model adapter files (in the *.safetensors format, as described by by321 (2022)). A detailed overview of the data structure can be found in our Supplementary Material, Figure S11.
2.2 Image-Level Analysis: Trends and Representations
To deepen our understanding of the visual material beyond initial impressions of the user-facing layer of the CivitAI platform, we examined all images and their metadata shared between January 2023 and December 2024. We first quantified temporal trends in the distribution of NSFW classification levels assigned by the CivitAI platform, as extracted from image metadata. We then analyzed the images themselves with regard to the demographic characteristics of depicted subjects, employing a classifier trained to infer age and gender to reveal broader trends in the generation of human figures.
2.2.1 Quantifying NSFW Trends
A previous study described by Palmini et al. (2024) revealed a disproportionate increase in NSFW imagery compared to SFW on CivitAI. The NSFW ratio, defined as the proportion of all images flagged as NSFW and assigned any browsing level other than 1 (which is SFW), increased from 56% in October 2023 to 73% in April 2024. Additionally, Wei et al. (2024) demonstrated that NSFW images receive more social engagement on CivitAI than SFW images. Our current study extends the dataset to include model metadata and covers a full two-year period, from January 2023 to December 2024. With this extended dataset, we confirm a continuation of the trend, with the NSFW ratio increasing from 41% in January 2023 to 80% by the end of December 2024. This trend is visualized in Figure 4, which presents a stacked bar diagram showing the color-coded proportions of over 40 million images according to their assigned NSFW browsing level, posted to CivitAI between January 2023 and December 2024. A more detailed account of monthly trends in image generation and NSFW classification is provided in Table S1 in the Supplementary Material.
Figure 3: Monthly distribution of generated images categorized by the NSFW browsing level assigned by CivitAI.
Figure 4: Inferred age and gender predictions using MiVOLO on a 0.1% sample of temporally representative images from the 40 million image dataset.
2.2.2 Demographic Patterns in Generated Images
To analyze and quantify the demographic distribution of human subjects depicted in generated images shared via CivitAI, we employed the MiVOLO (Multi-Input VOLO) model introduced by Kuprashevich and Tolstykh (2023). This is a transformer-based system designed to predict age and gender from images, even when faces are partially or fully occluded, while excluding images without human subjects. The results are shown in Figure 4, which shows the distribution of estimated gender and ages and the corresponding NSFW browsing levels.
The gender and age estimation was performed on a sampled subset of the original 40-million dataset. The subset consists of 40,636 images, representing approximately 0.1% of the full dataset. It was created by sampling, ensuring a temporally balanced and distributionally representative sample. Using the MiVOLO classifier, we identified 29,073 human subjects in 40,636 images, discarding images with no human subjects or a low confidence level for gender classification. The results of gender classification indicate a stark over-representation of female subjects, with an average female-to-male ratio of 6.24 to 1 across the dataset. NSFW co-occurrence rates were also notably higher for female subjects, with 85.44% of images depicting women with a browsing level other than 1, compared to male subjects, where the NSFW incidence was 73.66%. The average age for female subjects was 23.67 years, with a standard deviation of 5.1, while male subjects had a higher average age of 31.16 years with a standard deviation of 9.09 years. This indicates a narrower age range and a strong emphasis on sexualized portrayals of younger women, in contrast to a broader age distribution and more varied depictions of male individuals. A more detailed overview of these results, including a monthly breakdown of estimated gender and age distributions in sampled images, proportions across NSFW browsing levels, female-to-male ratios, and mean age statistics for each group is given in Table S2 in the Supplementary Material.
Several limitations should be noted in this analysis: CivitAI assigns browsing levels using Amazon Rekognition, unspecified open-source NSFW classifiers, and manual reviews. NSFW classifiers disproportionately misclassify images of female subjects as explicit compared to male subjects (Leu et al., 2024), potentially skewing NSFW label distributions towards higher-sensitivity categories for females. Additionally, MiVOLO’s binary gender classification relies on visually inferred traits, ignoring non-binary and gender-diverse identities, and may produce uneven error rates, especially for darker-skinned individuals (Buolamwini and Gebru, 2018). Age estimation models also introduce biases; for instance, digital ageism can distort the representation of older adults, particularly older women (Chu et al., 2022).
2.3 Model-Level Analysis: Themes and Biases
While generated images often attract the most attention, it is important to recognize that platforms like CivitAI primarily function as model-sharing hubs. The distribution and adaptation of models can reinforce ethically concerning trends, such as the hypersexualization and objectification of women, often more systematically than individual images. Accordingly, we shifted focus from shared images to the generative models behind them. Of the 232,164 assets published on CivitAI between November 2022 and January 2025, only 4.28% were standalone checkpoint models, while 94.04% were model adapters, which require a compatible checkpoint to generate images. A breakdown of asset types appears in Figure S12 in the Supplementary Material. To examine the thematic focus of these models, we began with an analysis of user-generated promotional tags. We then evaluated gender bias reproduction by running inference tests on the ten most popular checkpoint models. Next, we examined thirty popular model adapters, such as LoRA models, assessing their function in augmenting the visual repertoire of base models. Within this group, we identified a specific subset, referred to as deepfake adapters, which were designed to replicate particular individuals and made up about 15% of all adapters.
Figure 5: Co-occurrence network of the 300 most frequent promotional tags across all assets shared on CivitAI with the top three nodes indicating total occurrence rates and opacity and thickness intensity of co-occurrence rates.
2.3.1 Thematic Overview of Models shared on CivitAI
To understand dominant thematic orientations, we collected promotional tags added by model creators on CivitAI. These tags categorize and describe thematic, aesthetic, and functional attributes of models and adapters during the publishing process. We analyzed their frequency and co-occurrence across the platform. Figure 5 (interactive version available [^4] An interactive version of this graph is available here: https://pm-paper-viz.github.io/visualizations/figure_5.html) presents a co-occurrence network of the 300 most common tags. Each node represents a tag, with size indicating frequency. Edges indicate co-occurrence, with thicker, darker lines representing stronger connections. To aid visual interpretation, the size of nodes was normalized, and the top three most frequent nodes indicate their absolute occurrence counts. Anime aesthetics overall appeared to be a prevalent theme on the platform, with the tag “anime” appearing in roughly 30% of models. Additionally, tags denoting gender were highly prevalent. The tag “woman” appeared in 29.2% of shared models, “girls” in 15.7%, and “female” in 11.2%. Taken together, these figures indicate a pronounced emphasis on women and girls as depicted subjects across all assets exchanged on the platform. These tags were also commonly associated with tags such as “sexy”, “pose”, certain hairstyles and attire, and visual characteristics such as portraiture. The most active cluster revolved around character and anime and was connected to franchises like Dragon Ball, One Piece, and Gundam, and descriptors like TV show, visual novel, and game character. This cluster reflected the strong influence of fan culture and pop media. Among the various thematic patterns observed, the prominence of gendered tags stood out as one of the most striking results, highlighting deeper structural issues within the platform’s content ecosystem. Tags such as “woman” and “girl” not only dominated quantitatively but were also frequently linked to sexualized terms.
2.3.2 Bias in Model Checkpoints
To visualize what embedded discrimination in popular diffusion models means in terms of visual output, we performed an experiment on the top ten most popular checkpoints ranked by download count. Checkpoints are saved versions of a pretrained model, containing all necessary parameters to perform inference tasks without additional training and serving as the foundational assets for image synthesis, whereas adapters such as LoRA function as extensions merged into checkpoints to tweak their generative capabilities. To assess gender bias in popular checkpoint models shared on CivitAI, we performed a simple inference experiment based on comparing outputs of two different neutral text prompts. We selected the ten most popular checkpoint models according to download count available on the platform as of January 19th, 2025. For instance, Realistic Vision, the most downloaded asset at the time, had been downloaded over 1.5 million times, illustrating the scale of model adoption and its potential impact on generated content. We prompted each checkpoint model separately with the single-word prompts “woman” and “man”. Prompt embeddings were generated using the standard CLIP text encoder expected by each checkpoint. All images were generated using consistent parameters: CFG scale 6.5, Euler sampler (Qi et al., 2024), and 15 inference steps. To ensure comparability across outputs, we initialized the generation process with the same random noise seed for each model.
Figure 6: Top 10 most popular standalone models (type: checkpoint) as of 19th January 2024, prompted with the neutral terms woman and man using identical parameters. Initial noise seeds were iterated to ensure a representative, uncurated sample. Photorealistic and semi-realistic nudity was pixelized.
The selected checkpoints and their inference results are visualized in Figure 6. For comparison, we also included results from Stable Diffusion 1.5 and Stable Diffusion XL 1.0, since most of the selected checkpoints, (aside from Pony Diffusion), are derivatives of these models. The generated outputs across all models revealed a stark difference in gender representation: women were predominantly depicted nude, while men were frequently portrayed in professional attire.
While the baseline models, Stable Diffusion 1.5 and Stable Diffusion XL, still exhibit noticeable gender biases, these biases are even more pronounced in the model derivatives popular on CivitAI. It must be noted that our selection of checkpoint models was based on platform popularity metrics without restricting for model generality or application domain. As a result, the sample includes both general-purpose models and specialized models (e.g., anime-stylized, NSFW-oriented). This choice reflects the actual diversity of models encountered by end-users on CivitAI and allows us to capture platform-wide trends in gender representation. However, it also introduces specific variations that may intensify gender biases. Taken together, these findings point toward embedded discrimination within the platform’s model ecosystem, where user preferences, model fine-tuning practices, and training data biases converge to systematically reproduce and amplify narrow and hypersexualized portrayals of women.
2.3.3 Most Popular Model Adapters
The core aim of model personalization is to enable users to generate synthetic imagery with a degree of specificity and control that general-purpose models cannot achieve through prompting alone. This raises an important question: what kinds of “personalized visual worlds” are being created and shared in practice most commonly? To explore this, we examine the most popular model adapters on CivitAI. We identified the 30 most downloaded adapters on January 19th, along with their promotional tags, the compatible base architecture (base model), and download counts.
Figure 7: Top 30 most popular model adapters on CivitAI, including the number of downloads, adapter type, associated base model, and user-assigned promotional tags. Textual Inversion was abbreviated with “TI”.
A brief examination of the promotional tags of the most popular models confirms the thematic orientation identified in our broader tag analysis, with frequent references to tags such as “anime”, “female” and “sexy”, suggesting a strong emphasis on gendered and sexualized representations. Indeed, a closer look at the 30 most downloaded adapters reveals a strong presence of adapters implying pornographic themes. Many of the promotional tags accompanying these models specifically emphasize female subjects or body parts, with three models focused solely on the depiction of vaginas. Disconcertingly, one of the most popular adapters, Nudify XL: Better Bodies, directly references the practice of non-consensual “nudification” where photos of dressed individuals are manipulated to appear naked. Surprisingly, even models focused on stylistic variation often include tags like “female”. While CivitAI does not disclose how tags influence visibility, it is likely that popular tags improve discoverability. This can create a feedback loop: widely used terms like “female”, “girl”, or “sexy” not only reflect user interest but also shape platform norms, as creators adopt them to boost exposure.
2.3.4 Deepfake Adapters and Targeted Subjects
One of the most ethically concerning applications of model personalization is the creation of model adapters intended to replicate real individuals. To better understand the scope and impact of this phenomenon, we focused on analyzing the subset of such adapters, investigating their prevalence, common use cases, and the demographic characteristics of the individuals they target. According to Deeptrace’s 2019 report, 96% of all deepfake content online was pornographic (Ajder et al., 2019), a trend that intensified by 2023 when a report by Home Security Heroes found that 98% of deepfake videos were pornographic and 99% of the victims were women (Home Security Heroes, 2023; Birrer and Just, 2024).
Although CivitAI prohibits users from publicly sharing explicit content involving real people, it has no control over how models are used once downloaded or what is shared on external platforms. This creates a significant blind spot in the enforcement of ethical guidelines, as deepfake models can still be deployed in harmful ways beyond the platform’s visibility. Given this regulatory gap and the continuing risk of abuse, we sought to understand who is being targeted by these models. To gain more insight into the individuals currently targeted by the practice of non-consensual deepfakes created from models shared on CivitAI, we annotated our dataset of deepfake adapters using a large language model, similarly to the approach pioneered by Wei et al. (2024). This allowed us to identify the most commonly associated professions and countries linked to those individuals.
(a) Sunburst chart showing the distribution of individuals segmented by gender (inner ring) and professional category (outer ring).
(b) Sunburst chart showing the distribution of individuals segmented by associated country (inner ring) and professional category (outer ring).
Figure 8: Sunburst charts depicting the distribution of 16,078 individuals targeted by deepfake models available on CivitAI, categorized by LLM-predicted profession, gender, and associated country.
In our data set of models and adapters, we found 15.27%, 33,804 model adapters were tagged “person of interest” as discussed in Section 3.1, which means that these were tailored to replicate identifiable individuals. The individual’s name is rarely included in the model tags and is typically indicated in the model name. However, many deepfake model names employ obfuscation techniques such as leet speak, where letters are substituted with visually similar numbers or symbols, in order to evade detection or moderation. To find more data on the individuals targeted by deepfake models, we standardized model filenames by converting obfuscated text into regular characters and removing version numbers and emojis. We then used the en_core_web_trf model (Explosion, 2021) from spaCy 3 (Montani and Honnibal, 2021) for named entity recognition (NER), and subsequently queried DeepSeek (DeepSeek-AI et al., 2025) to infer likely professions and countries, based on the NER-identified name and contextual hints included in the model’s promotional tags. For each entry, we retrieved up to three ranked professions and chose the top-ranked item. Entries marked as uncertain were labeled “unknown” and excluded, as were those referring to fictional characters or trained on multiple individuals, such as e.g. an adapter named “bond girls”. After filtering and aggregating models associated with the same individual, 16,078 unique instances of individual names remained.
To visualize our results, we use sunburst charts, which represent nested categories as concentric rings, to illustrate how inferred professions are distributed across gender and country associations. Figure 8 includes two sunburst charts that visualize the distribution of inferred professions among individuals represented in the dataset. Figure 8(a) shows this distribution aggregated by gender. While Figure 8(b) breaks down the same professional categories by the most frequently associated countries.
Our results indicate that actors emerged as the most common profession across both genders and all countries, with the exception of Japan. Among women, adult performers were the second most frequent category, whereas public figures ranked second among men.
The majority of targeted individuals were associated with countries in the Global North, specifically the United States, the United Kingdom, and Japan. Cultural factors appear to influence who is targeted: for example, Japan showed a notably high presence of voice actors compared to other countries, likely reflecting the country’s strong anime and gaming industries. Sports professionals were also frequently identified, a pattern that may be attributed to the increased media exposure surrounding international events such as the 2024 Olympics.
This analysis offers insight into the cultural and professional context of deepfake victims but is limited by simplifying assumptions and methodological constraints. Language models may misclassify individuals or overlook multiple roles, and restricting profession predictions to three options while using only the first simplifies complex identities. DeepSeek can conflate distinct geographic regions. Nonetheless, the analysis provides a structured entry point for exploring the social dynamics behind deepfake creation.
The creation of deepfake explicit images is a clear abuse that highlights the risks of model personalization. Yet beyond these malicious uses lies a deeper, systemic problem: the spread of misogynistic and ethically troubling content through the very structures of model training. These issues often arise not from intent, but from seemingly neutral processes like image-text pairing. In the next section, we examine how such mechanisms embed gendered and sexualized assumptions into personalized models.
2.4 System-Level Analysis: Configurations and Training Practices
Even when personalization workflows do not explicitly aim to sexualize or objectify individuals, the training mechanisms themselves can introduce and reinforce such representations. Training model adapters requires training data in the form of image-caption pairs. Manually generating captions for each image is time-consuming so auto-tagging systems are commonly used to automate this process by generating descriptive tags or image captions based on the visual content of each image. While many open-source TTI models, tools, and communities exist, only a few auto-captioning systems are widely adopted.
One of the commonly employed auto-tagging system is CLIP Interrogator, which generates captions by calculating cosine similarities between image features and a structured vocabulary using the CLIP model (Radford et al., 2021). Its vocabulary is organized into categories such as artists, mediums and art movements (pharmapsychotic, 2022). While a few terms reference violence, the vocabulary does not include explicit sexual terms.
The other common approach involves Danbooru-based auto-tagging systems. Danbooru [^5] Danbooru https://danbooru.donmai.us/, (Accessed: 20 November 2025) is a collaboratively tagged, anime-focused imageboard, whose vocabulary includes a high prevalence of sexually explicit terms, while offering comparatively few labels for domains such as nature or art history (Britt, 2019). Several tagging systems have been trained on the Danbooru dataset, the most prominent of which is DeepDanbooru  (Kim, 2024).
Danbooru-based tagging systems are commonly integrated into popular tools for training LoRA adapters. For instance, the image generation and model training GUI apps, kohya-ss (bmaltais, 2025) and Stable Diffusion WebUI, support both Danbooru-based systems and CLIP Interrogator in their model training pipelines, allowing users to choose between them. In the well-known and widely adopted open-source TTI tool ComfyUI, there is no default tagging system, but users can install extensions such as WD-14 Tagger (smilingwolf, 2022) which is a Danbooru-based auto-tagging system.
Although the Danbooru dataset contains only anime-style images, tagging systems trained on it are frequently applied to photorealistic datasets. The Danbooru vocabulary is shaped by the norms of the Danbooru imageboard’s user base. While some material may be objectionable, it can still be protected as artistic expression within stylized contexts such as anime. However, the same content might cross legal boundaries if rendered in a photorealistic style. Captioning systems trained on the Danbooru repository typically produce short, discrete tag sequences, such as for example “1girl, red dress, white background”. In contrast, CLIP Interrogator generates full-sentence descriptions that resemble natural language, for example “a woman wearing a red dress standing in front of a white background”. Because these systems produce characteristically different outputs, it is possible to infer which captioning system was used by identifying patterns in the metadata. This approach is particularly effective for Danbooru-based systems, which employ highly specific and recognizable tags such as “1girl” or “1boy”. To categorize model adapter training keywords and better understand the overall semantic structure of the data, including the presence of potentially problematic or sensitive content, we created a structured taxonomy of Danbooru tags derived from the publicly available tag groups on the Danbooru imageboard. The resulting hierarchy, which contains 35,559 unique terms together with their categories, aliases, and implications, is shown in Figure S13 in the Supplementary Material.
To understand the impact and potential problematic uses of auto-tagging systems in the context of CivitAI, we explore which tagging systems are most commonly used in model adapters shared on the platform. We analyse a subset of the 40,000 most downloaded CivitAI assets. Within that subset, we can identify the auto-tagging systems used for model training for approximately 11,000 models by analyzing the available training data. More specifically, we analyze the image captions preserved in model metadata fields, which provide insight into the types of language used to describe training content.
To identify the source of captions in training metadata, we compared each set of tags to the characteristic vocabularies and structural patterns of both CLIP Interrogator and Danbooru-based systems. If fewer than five matches were found, or if tags or metadata were missing, the captioning system was classified as unknown.
Our results indicate that only 14% of the model adapters in our sample were trained with captions matching CLIP Interrogator patterns, while more than 70% were matching Danbooru-based tagging systems. This trend raises ethical concerns, particularly within the subset of deepfake adapters, where 42.6% of those flagged as representing real individuals included Danbooru-derived tags.
Figure 9: Subdivision of the 40,000 most downloaded CivitAI assets by type, presence of extractable training data, and occurrence of problematic tags.
To illustrate the overall landscape of model personalization and the influence of potentially problematic training tags, Figure 9 presents a Sankey diagram that categorizes the 40,000 most-downloaded models on CivitAI, indicating how many are adapters, how many include extractable textual training data, and within that subset, how many captions contain explicit keywords. Among these, we further highlight the number of models that contain specifically abusive terms.
Within the group of models for which textual training data were available, about half contained at least one tag mapped to the Danbooru “sex” category, based on the Danbooru taxonomy presented in the Supplementary Material in Figure S13 and were therefore marked as “explicit”. Roughly 5.6% included the keywords “loli” and/or “shota”, terms commonly used to denote sexualized depictions of pre-pubescent girls and boys. About 2.1% included the keyword “rape”. These results underline how the choice of captioning tools and training configurations shapes the personalization of TTI models. Our findings suggest that these systems not only automate annotation but also embed norms originating from specific niche internet communities, which normalize sexualized depictions of minors and reinforce misogynistic tropes.
When combined with broader trends in user behavior and platform features, such as those observed on CivitAI, these practices contribute to the widespread reproduction of narrow aesthetic and thematic patterns. In the following discussion, we contextualize these findings within the larger context of model personalization, drawing connections across the preceding sections to examine the social, technical, and ethical implications of current personalization workflows.
3 Sociotechnical Implications and Pathways for Intervention
The following discussion is organized around four principal thematic domains identified through our analysis. First, we assess the systemic dynamics of platform decay, wherein engagement-driven algorithms and monetization strategies appear to incentivize exploitative feedback loops that undermine community integrity. Second, we analyze the entrenchment of gendered biases through the lens of perpetual misogyny, highlighting how historical patterns of objectification are rearticulated and amplified within generative AI outputs. Third, we focus on the proliferation of deepfakes and the ethical implications of the increasing accessibility of tools that enable the non-consensual replication of personal likenesses. Finally, we explore potential modes of intervention for mitigating harm, including technological inventions, regulatory measures and community governance practices.
3.1 Platform Decay
What began as a promising creative breakthrough in TTI generation and model personalization, has devolved into a pipeline for the large-scale production of sensational, biased, and abusive content. The open-source nature of TTI technologies, proclaimed as a democratizing force in generative AI, has also enabled the propagation of models that perpetuate hypersexualized imagery and non-consensual deepfakes. This shift reflects a broader structural trend described by the phenomenon of enshittification, as coined by Doctorow (2025), also known as platform decay. It describes the process by which a platform originally offering value to its customers grows increasingly exploitative under the imperative to recoup initial investments. Several indicators suggest a descent into a self-reinforcing feedback loop of platform decay. These include a dramatic increase in NSFW imagery, from 41% to 80% in two years, as well as the community’s normalization of deepfakes, misogynistic tropes, and other exploitative content.
While moderation and community health measures typically focus on a platform’s internal dynamics, toxicity within internet subcultures often extends beyond these boundaries, affecting the broader social context (Trott et al., 2022). Social media entities are explicitly marketed as platforms rather than publishers, a distinction that suggests they are not bound by the ethical responsibilities typically expected of traditional publishers (Applebaum, 2024). In the case of model-sharing hubs, the consequences of this distinction may be especially problematic, as the output of models and model adapters, by design, serves as a means to generate and perpetuate content across the internet.
As the commercialization of open-source TTI has prioritized profit over inclusivity, model-sharing hubs tend to favor sensational content to boost engagement metrics, which may exert a lasting negative impact on diversity and equitable representation in AI-generated media. Open-source technologies originally developed within collaborative, community-driven contexts, often transition into consumer products and are absorbed into profit-oriented ecosystems, as demonstrated by the ubiquity of Linux-derived operating systems and the widespread adoption of Meta’s PyTorch framework (Widder et al., 2023). Similarly, TTI models can be regarded as open-source software artifacts. As such, these models bear the potential to cause extensive downstream harm, particularly when biased, hypersexualized, or non-consensual content is widely replicated and disseminated without meaningful oversight or accountability.
3.2 Perpetual Misogyny
According to Manne (2017), misogyny is not merely the hatred of women or the ideological framing of women and girls as inferior, which constitutes sexism. Rather, it is performative: a hostile set of mechanisms and actions designed to reinforce patriarchal power structures. The over-sexualization and objectification observed in our analysis exemplify such mechanisms, as they perpetuate harmful stereotypes and contribute to a culture that marginalizes women and may deter their participation on the platform.
Drawing on feminist and psychoanalytic critiques, from Mulvey’s notion of the “male gaze” (Mulvey, 1989) and Doane’s readings of female representation (Doane, 1991) to Berger’s critique of Western art (Berger, 1972), and recent scholarship summarized in (King, 2024), our work demonstrates that digital and AI-generated imagery echoes the historical objectification of women’s bodies. Visual culture frequently compartmentalizes the female body, as seen in advertizing, which may explain why we cognitively process images of women similarly to how we recognize sets of objects, whereas we tend to perceive images of men as whole entities (Gervais et al., 2012). Our analysis of the most popular model adapters indicates the persistence of this trend in the context of generative AI, where women are frequently depicted through fragmented, sexualized representations that emphasize specific body parts. The reduction of women to isolated body parts through objectification extends into fantasies of replacement, where the human subject is erased and substituted with artificial representations devoid of autonomy or consent.
Erscoi et al. (2023) demonstrate how the concept of Pygmalion Displacement, derived from the myth of the Cypriot king who preferred the marble statue Galatea over real women, remains a consistent trope in science fiction as well as in recent technological advancements. The sexual fantasy of the woman automaton, a machine created for male sexual gratification and/or menial tasks, continues to shape representations of artificial femininity. In the context of generative AI, this fantasy is not only perpetuated but also operationalized through the personalization of models that allow users to generate synthetic female figures tailored to idealized and often hypersexualized specifications. Often, these artificial women are created by men and for male consumption, by means of a technology that is the distillate of billions of images, among them hundreds of thousands of women’s bodies with partly questionable origin (Birhane et al., 2021), fine-tuned with data often used without the consent of the individual represented. Historically, Pygmalion fantasies coincide with the systematic displacement of women from technological fields (Abbate, 2012), a trend mirrored in the evolution of open-source TTI from its originally more diverse, community-driven origins, documented by Oppenlaender (2022), to a space increasingly commercialized and dominated by male interests.
3.3 The Proliferation of Deepfakes
As problematic as objectification and Pygmalion fantasies may be, they differ fundamentally from the disruptive impact that ultra-realistic deepfakes can exert on an individual’s life. The consequences of sharing pornographic deepfakes can, depending on the circumstances or cultural context of the victim, range from a minor nuisance to suicide (Gavrilović Nilsson et al., 2019) or homicide, specifically femicide (Ng, 2025). While this is tragic on an individual level, the now readily available technology may have detrimental effects on society as a whole. The ability to manipulate someone’s likeness, face, and/or voice is not only a severe violation of that individual’s dignity, it undermines the foundational principles of liberal democracies, including autonomy, consent, trust-based exchange and the right to self-ownership.
While platforms such as CivitAI are not responsible for the mere existence of such technologies, they actively normalize and even incentivize their creation. In the case of CivitAI, this occurs through a “celebrity-creator leaderboard”. This term is itself a misnomer, as it extends to individuals without a significant public visual record, such as YouTubers, Instagrammers, many of whom might not have the financial means to remove widespread AI-generated non-consensual pornography from the internet. As we discussed in Section 3.4, our survey of 221,304 CivitAI assets found that 15.27% (33,804 model adapters) were flagged as person of interest, meaning they are tailored to replicate the likeness of real individuals. Although CivitAI currently does not permit the sharing of pornographic images containing real individuals, they do not control the local use of models once downloaded. Reacting to pressure from payment processors, CivitAI recently introduced limited policy adjustments, such as removing the “celebrity” tag from certain models, restricting the sharing of incest-related pornography, and adjusting their feed organization so that imagery of children or real people may not be displayed between pornographic content any longer (Anderson, 2025). However, to this day, all deepfake adapters remain available on the platform. Individuals finding themselves digitally immortalized against their will on CivitAI may appeal for the removal of adapters through a formal appeal process, shifting the burden of action onto the victim, most likely after the model has already been downloaded multiple times and potentially redistributed across other platforms.
3.4 Pathways to Harm Mitigation
In line with constructivist perspectives on technology (Hayles, 2025), we argue that technological development, including generative AI, is not determined but contingent, and as such, subject to complex sociotechnical dynamics. All technology is shaped by the value systems from which it emerges, stakeholder decisions, and underlying power structures (O’Connor and Liu, 2024). The LAION datasets, foundational to many TTI models despite being “not recommended” for commercial use (Buschek and Thorp, 2025), have been called the “original sin” of TTI (Salvaggio, 2024) due to their harmful content (Birhane et al., 2021, 2023; Thiel et al., 2023). While model personalization could have mitigated inherited biases, our study finds that public participation has instead deepened LAION’s toxic dynamics, both systemically and socially through personalization hubs. These hubs not only amplify technical flaws but also expose deeper societal pathologies, confronting us with visual symptoms of racism, ageism, misogyny, and dehumanization. Such manifestations may spark critical discourse on the structural roots of these harms.
While recognizing these outputs as symptoms of deeper structural problems is important, curbing immediate harms through safeguards at the platform level remains essential. This requires innovative responses from platform operators, particularly in developing nuanced content generation mechanisms. For instance, Haidra [^6] https://haidra.net/ (Accessed: 11 November 2024), a decentralized compute provider, developed a safety mechanism, that makes use of DeepDanbooru’s knowledge of problematic concepts, to block abusive image generations on their hardware (tazlin, 2023).
Despite such innovative solutions, some issues require more immediate and explicit policy interventions, especially the growing threat of deepfakes. For example, the model-sharing hub Hugging Face restricts the sharing of all models that mimic the likeness of real people, unless they are shared with the explicit consent of the individual depicted (Hugging Face, 2025). Yet, even when model-sharing hubs implement policies, significant harm may already have spread beyond their control. Individuals affected by digital violence require both retroactive and preventative support. StopNCII [^7] https://stopncii.org/ (Accessed: 10 November 2024) offers preventative tools that allow individuals to hash non-consensual intimate imagery (NCII) so partnering online content and social media platforms can then use these hashes to detect and block such content. Amiinporn [^8] https://amiinporn.io/ (Accessed: 20 April 2025) by the NGO Digital Dignity, supports victims by helping detect and remove existing non-consensual content on sites that provide database access. Additionally, FaceSafe [^9] https://facesafe.io/ (Accessed: 20 April 2025), represents a proactive approach aimed at preventing deepfakes and identity theft. Once launched, this subscription-based service will allow individuals to register a vector derived from extracted facial features in order to be notified if their likeness appears on any participating platform. In addition, these issues are being tackled on the technological level through the development of various protective techniques, methods designed to prevent misuse or reproduction of sensitive content. For example, image perturbation methods like Glaze (Shan et al., 2025), Nightshade (Shan et al., 2024), and Mist (Liang and Wu, 2023) introduce small perturbations that confuse earlier diffusion models, exemplified by Stable Diffusion 1.5. However, these defenses tend to lose effectiveness against newer architectures like FLUX and Stable Diffusion 3.5, which employ more advanced designs, including transformer-based components, enhancing their resilience to visual noise. As highlighted by Zhao et al. (2024), most traditional perturbation methods fail against fine-tuning techniques like LoRA, which adapt specific attention layers within these models. Newer techniques like MetaCloak (Liu et al., 2024) have been developed to be specifically effective against such fine-tuning methods. However, due to the rapid development of models, image perturbation techniques must be constantly updated. As a potential protective mechanism, these methods could be integrated into social media platforms as experimental features, for example, as an optional image filter applied upon upload. But the effectiveness of these protective mechanisms ultimately depends on the willingness of social media and content creation platforms to protect their users from non-consensual imagery, a commitment that is currently lacking (Li et al., 2024).
Conclusion
Patterns of structural discrimination inadvertently introduced into large-scale foundation TTI models are likely to become overshadowed by the biased models and adapters deliberately infused into the digital ecosystem through open-source model-sharing hubs and their users. As shown in this study, on model-sharing hubs like CivitAI biases, particularly those reinforcing hypersexualization and gendered objectification, are amplified and normalized. The rapid rise of NSFW content, the over-representation of young female subjects, and the prioritization of sensational content to drive engagement reflect an exploitative, even abusive dynamic. Additionally, structural discrimination embedded in today’s open-source TTI tools and models have the potential to cause significant downstream harm as they might become widely adopted and even integrated into future consumer applications.
Even more alarming is the normalization and incentivization of deepfake model adapters, which fundamentally violate individual dignity, autonomy, and the core democratic principles of self-ownership and consent. While technological development is not deterministic but contingent, the current trajectory highlights how platform dynamics can exacerbate social pathologies rather than foster equitable access. Yet the contingent nature of technology, shaped by online communities, platform operators, lawmakers, and society as a whole, also creates opportunities for intervention. Model-sharing hubs and social media platforms both have the capacity to implement safeguards that can limit the spread of abusive practices such as deepfake creation and abusive imagery. Strengthening responsible moderation, supporting ethical model development, introducing technical safeguards, and collaborating with NGOs focused on preventing digital violence, are necessary steps to adjust the current trajectory. Rather than reacting after harm has occurred, all stakeholders have the opportunity to innovate proactively and to collectively shape the future direction of model sharing and personalization ecosystems.
Acknowledgments
We are grateful to Amanda Wasielewski, Tristan Weddigen, Maria-Teresa de Rosa Palmini, Tania de León Yong for their valuable feedback and suggestions on this paper. Special thanks to Lasse Scherffig and Thomas Hawranke for their guidance in shaping the initial proposal for this study. In addition, we thank Jonas Schnabel and Yannick Schuchmann of Digital Dignity e.V. for their insightful input.
Funding
This work was supported by Swiss National Science Foundation (Ambizione Grant 216104)
Supplementary Material
.5 Supplementary Figures for Section 3.1: Data Collection and Categorization
1{ 2 "id": 40603651, 3 "nsfw": false, 4 "browsingLevel": 1, 5 "createdAt": "2024-11-17T12...738Z" 6 ... 7 "stats": { 8 "likeCount": 5, 9 }, 10 "meta": { 11 "prompt": polka-dotted squirrel, 12 ultra wide angle::3 BREAK 13 "civitaiResources": [ 14 { 15 "type": "checkpoint", 16 "modelVersionId": 691639, 17 }, 18 type: "lora" 19 ] 20 }, 21 "username": "9gc90", 22 "baseModel": "Flux.1 D" 23}
(a) Exemplary image metadata (lines omitted) containing the text prompt, as well as parameters and models and adapters used in the image genesis.
1{ 2 "id": 600987, 3 "name": "Naive Abstract", 4 "type": "LORA", 5 "poi": false, 6 "nsfwLevel": 1, 7 "stats": { 8 "downloadCount": 136, 9 }, 10 "creator": { 11 "username": "9098f7f" 12 }, 13 "tags": [ 14 "abstract" 15 ], 16 "modelVersions": [ 17 { 18 "publishedAt": "2024-11-2TZ", 19 "hashes": { 20 } 21 } 22 ] 23}
(b) Exemplary metadata of an model adapter e.g. LoRA, containing a unique identifier, publishing date, person-of-interest flag etc.
Figure S10: Addendum to Section 3.1 Typical image and model metadata as obtained via REST API from CivitAI. Some lines have been omitted for clarity. All assets contain unique identifiers assigned by CivitAI, as well as creation dates, usernames etc.
1"metadata":{ 2 "ss_max_train_steps":"3940", 3 "ss_tag_frequency":{ 4 "an image of a woman...":16, 5 "a white face":6, 6 "adding depth ...":3, 7 "creating contrast ...":2, 8 "adding a pop of color ...":2, 9 "the background is ...":2, 10 "an image of a woman ... 11 she is wearing a black corset... 12 she has a black glove ... 13 her eyes are blue and her ... 14 in a ponytail and shes ... 15 background is brick wall...":2 16 }, 17 "ss_caption_tag_dropout_rate":"0.0" 18}
(a) Example of metadata containing textual training data typical for image captions generated using the CLIP Interrogator captioning system
1"metadata":{ 2 "ss_num_train_images":"108", 3 "ss_tag_frequency":{ 4 "kasumi":216, 5 "1girl":108, 6 "solo":108, 7 "bare shoulders":84, 8 "japanese clothes":79, 9 "blunt bangs":63, 10 "breasts":62, 11 "weapon":53, 12 "hair ornament":51, 13 "sword":51, 14 "kimono":49, 15 "pink kimono":45 16 }, 17 "model_name":"EMS-263145.safetensors" 18}
(b) Example of metadata with textual training data that correspond to the pattern typical for captions generated using systems built involving Danbooru tags.
Figure S11: Addendum to Section 3.3.4 Comparison of metadata extracted from downloaded model adapters’ *.safetensors files, showcasing differences between captioning training data with CLIP Interrogator compared to Danbooru-based image captioning systems.
.6 Supplementary Figures for Section 3.2: Understanding Image Contexts
Table S1: Monthly statistics from the Civiverse-40M dataset, covering November 2022 to December 2024. The table reports the total number of generated images, the number classified as NSFW (Not Safe for Work) and non-NSFW, the average number of images generated per day, and the NSFW ratio (the proportion of NSFW images relative to the total image count).
Year Month Total Images NSFW True NSFW False Avg. Daily Images NSFW Ratio
2022 11 3451 419 3032 115.03 0.121
2022 12 5976 1529 4447 192.77 0.256
2023 1 26789 11051 15738 864.16 0.413
2023 2 65916 33907 32009 2354.14 0.514
2023 3 133000 67968 65032 4290.32 0.511
2023 4 164187 82499 81688 5472.90 0.502
2023 5 230673 115468 115205 7441.06 0.501
2023 6 251083 126848 124235 8369.43 0.505
2023 7 324691 162005 162686 10473.90 0.499
2023 8 334493 151302 183191 10790.10 0.452
2023 9 339205 176758 162447 11306.83 0.521
2023 10 388600 215200 173400 12535.48 0.554
2023 11 552638 322122 230516 18421.27 0.583
2023 12 661746 370372 291374 21346.65 0.560
2024 1 817261 478818 338443 26363.26 0.586
2024 2 926644 567173 359471 31953.24 0.612
2024 3 1328061 863795 464266 42840.68 0.650
2024 4 1854660 1352812 501848 61822.00 0.729
2024 5 2436790 1857087 579703 78606.13 0.762
2024 6 2968333 2358802 609531 98944.43 0.795
2024 7 3518577 2854555 664022 113502.48 0.811
2024 8 3958639 3110646 847993 127698.03 0.786
2024 9 4156712 3289692 867020 138557.07 0.791
2024 10 4559842 3680924 878918 147091.68 0.807
2024 11 5082820 4094030 988790 169427.33 0.805
2024 12 5539773 4405664 1134109 178702.35 0.795
N Images: 40,630,560     (NSFW: 30,754,835)
Notes: All images were posted publicly to CivitAI between November 2022 and December 2024. NSFW-true images are those with Browsing Levels 2, 4, 8, or 16. NSFW-false images have Browsing Level 1. The NSFW ratio is calculated as the proportion of NSFW-true images out of the total monthly count.
Table S2: Addendum to Section 3.2, MiVOLO Age and Gender Estimations for 2023–2024
[…]
L1 L2–16 L1 L2–16
2023-01 26 48 9 38.46 61.54 75.00 25.00 3.25 22.26 (2.38) 30.93 (4.46)
2023-02 67 68 16 38.30 61.70 100.00 0.00 11.75 23.53 (4.95) 34.60 (6.52)
2023-03 133 117 30 37.36 62.64 45.45 54.55 8.27 23.94 (5.00) 32.40 (7.02)
2023-04 164 134 37 37.84 62.16 60.00 40.00 7.40 24.25 (6.90) 33.90 (9.99)
2023-05 229 242 52 35.53 64.47 70.83 29.17 6.33 24.40 (5.61) 32.24 (8.88)
2023-06 251 233 72 40.65 59.35 53.33 46.67 5.17 22.55 (4.41) 30.52 (9.03)
2023-07 317 271 75 39.30 60.70 64.10 35.90 5.15 22.81 (4.65) 32.16 (9.92)
2023-08 326 299 94 39.59 60.41 77.78 22.22 5.47 23.67 (6.60) 37.69 (8.54)
2023-09 316 287 58 30.53 69.47 74.29 25.71 6.46 23.73 (5.30) 30.46 (11.91)
2023-10 351 294 78 36.48 63.52 68.42 31.58 6.13 23.33 (6.33) 34.04 (9.39)
2023-11 560 474 134 27.79 72.21 41.67 58.33 6.12 23.13 (4.94) 34.84 (12.05)
2023-12 674 580 143 35.97 64.03 63.77 36.23 6.77 23.21 (4.88) 33.06 (11.14)
2024-01 752 659 155 31.49 68.51 52.22 47.78 5.61 23.32 (5.50) 33.53 (8.22)
2024-02 868 728 211 29.37 70.63 55.14 44.86 5.19 23.77 (5.65) 32.55 (8.44)
2024-03 1305 1094 358 24.61 75.39 46.21 53.79 6.31 24.11 (5.58) 32.42 (10.63)
2024-04 1733 1411 516 20.19 79.81 37.64 62.36 5.96 24.17 (4.96) 32.06 (9.42)
2024-05 2304 1769 704 17.11 82.89 25.57 74.43 6.54 24.14 (4.80) 30.84 (9.42)
2024-06 2857 2049 981 13.89 86.11 23.93 76.07 5.89 23.77 (4.79) 28.88 (8.36)
2024-07 3384 2416 1155 14.47 85.53 21.65 78.35 6.77 24.01 (4.96) 30.18 (9.71)
2024-08 3808 2814 1299 15.70 84.30 28.35 71.65 5.66 24.03 (4.90) 30.53 (9.11)
2024-09 3912 2840 1350 14.16 85.84 27.90 72.10 6.20 23.92 (4.54) 31.51 (8.64)
2024-10 4250 3224 1381 13.52 86.48 26.45 73.55 5.85 23.94 (4.79) 31.19 (9.83)
2024-11 4729 3458 1587 15.48 84.52 23.92 76.08 5.63 24.08 (5.24) 30.57 (8.82)
2024-12 5139 3564 1860 14.00 86.00 24.90 75.10 5.80 24.04 (4.83) 30.72 (8.61)
Addendum to Section 3.2: A 0.1% subsample ($n=40{,}636$) of the entire image dataset was used to estimate the demography of depicted subjects. MiVOLO was used to estimate perceived age and gender, from which we calculated the female-to-male ratio and average ages (with standard deviation) of female and male subjects. Note: person count may exceed image count as multiple persons may appear per image.
[…]
.7 Supplementary Figures for Section 3.3: Profiling Shared Models and Adapters on the Platform
Figure S12: Distribution of asset types shared on CivitAI, excluding generated images. Asset types include standalone models (referred to as checkpoints), model adapters (such as LoRA, LoCon, and Textual Inversion), and other resources such as workflows for ComfyUI, pose depth maps, or prompt “Wildcards” shared on the platform. Wildcards refer to placeholder tokens (e.g., early 2000’s flash photography) used to randomly construct prompt variations for diverse image generation. Pose depth maps are structural guides—such as body poses or 3D depth data—used by ControlNets to constrain and compose images.
.8 Supplementary Figures for Section 3.4: Insights from Training Configurations and Practices
Figure S13: Danbooru taxonomy of 35,559 unique Danbooru tags as extracted from the tag categories and assigned the tag hierarchy from the danbooru.donmai.us tag wiki.
Figure S14: Co-occurrence network map of textual training data based on discrete tags, as described in Section 3.4. The data was extracted from the model’s metadata files (*.safetensors) and visualized as a network graph. Each node represents a tag, and edges indicate co-occurrence relationships between tags, with line thickness and proximity reflecting the strength of their association. The tags are color-coded according to the Danbooru taxonomy, providing a visual categorization of semantic clusters and thematic groupings within the dataset. An interactive version of this graph is available at: https://pm-paper-viz.github.io/visualizations/figure_15.html.